(2x-5)(3x+2)=2(3x-11) what are we going to do with this expresion?
Answer by checkley79(3341)

(2x-5)(3x+2)=2(3x-11)
6X^2-15X+4X-10=6X-22
6X^2-11X-6X-10+22=0
6X^2-17X+12=0
(3X-4)(2X-3)=0
3X-4=0
3X=4
X=4/3 OR 1 1/3 ANS.
2X-3=0
2X=3
X=3/2
X=1.5 ANS.